OpenStudy (anonymous):

Rewrite the expression as a trigonometric function of a single angle measure.

OpenStudy (anonymous):

\[\frac{ \tan 4\theta +\tan 2\theta }{ 1-\tan 4\theta+\tan 2 }\]

OpenStudy (anonymous):

Theres suppose to be a \[\theta \] at the end of the '2'